# Avoiding matrix exponentials for large transition rate matrices

### 09 Dec 2023 v1

@@ 2.2.1 Method #1 – Trajectory-based MJP method (T-MJP)
When enough alternating samples are drawn, we obtain a model of the dynamics (that is, samples of $\theta$ informed by $\bar{s}$) without ever computing the matrix exponential itself.
@@ 2.2.2 Method #2 – Jump-based MJP method (J-MJP)
As with the previous method, for a sufficiently large number of alternate samples, we can avoid the matrix exponential entirely and still obtain the appropriate probabilities for the states given the kinetic parameter, $\theta$, and the number of virtual jumps for respective observations, $\bar{k}$. In Sec. 3.2.2 we will discuss practical challenges this poses.
@@
Rather than the complete set of $\bar{\rho}(0)\mathbf{B}_{\theta}^{k}$ required by T-MJP, algorithm 1 only requires saving the vector $\bar{\rho}\mathbf{B}_{\theta}^{k}$, denoted in the pseudocode as $\bar{\rho}^{*}$. Its outer loop is run $\max\limits_{i}k^{i}$ times. Given that $k^{i}$ is Poisson distributed at rate $T^{i}\Omega_{\theta}$, see Sec. 2.1.1, and $\Omega_{\theta}$ scales with $\max\limits_{n}{|a_{nn}|}$, the number of outer loop calls scales with $O(T\max\limits_{n}{|a_{nn}|})$, with $T$ being the largest collection time $T=\max\limits_{i}T^{i}$. Meanwhile, the most computationally intensive operation within each loop iteration is multiplying $\bar{\rho}$ by $\mathbf{B}_{\theta}$, which scales in time as $O([\mathbf{B}_{\theta}])=O([\mathbf{A}_{\theta}])$. Therefore the total time scaling of the state probability computation is $O(T\max\limits_{n}{|a_{nn}|}[\mathbf{A}_{\theta}])$. Under conditions a) and b) mentioned in the introduction, this is equivalent to $O(N^{2})$.
@@ 2.2.3 Method #3 – Remarginalized MJP method (R-MJP)
where $\varepsilon_{k^{\ast}}=1-\sum_{0}^{k^{\ast}}\ \text{Poisson}(k|\Omega_{\theta}t)$, is the upper bound in the approximation error. Above we choose to write a generic $t$ to represent the fact that (16) is valid for all times, but in practice the goal is to calculate the probability vectors at the collection times $T^{i}$. For practical reasons, we choose $k^{\ast}=\Omega_{\theta}t+6\sqrt{\Omega_{\theta}t}$, guaranteeing a modest error bound $\varepsilon_{k^{\ast}}<10^{-5}$ for $\Omega_{\theta}t>10$. While implementing the approximation in (16) is straightforward, we provide pseudocode detailing how to effectively calculate the summation in (16) in Algorithm 2.
@@
In order to calculate the probability vector at the observation time $\bar{\rho}(t)$ from $\bar{\rho}(0)$, the outer loop in Algorithm 2 is set to run $k^{\ast}=\Omega_{\theta}t+6\sqrt{\Omega_{\theta}t}$ times. Similarly to the previous method, this scales as $O(T\max_{n}|a_{nn}|)$. Within this loop, the most computationally expensive task is multiplication of an array of $N$ elements by $\mathbf{B}_{\theta}$, with cost scaling $O([\mathbf{B}_{\theta}])$. Consequently, the overall time cost scaling is $O(\max_{n}|a_{nn}|[\mathbf{B}_{\theta}])$. Again, when conditions a) and b) in Sec. 2.1.2 are met, this overall cost becomes $O(N^{2})$, while the Algorithm 2 requires only saving a single array of size $N$ besides the matrix $\mathbf{B}_{\theta}$, giving memory scaling $O(N)$.
@@ 2.2.4 Method #4 – Runge-Kutta
The most conceptually straightforward technique to avoid the matrix exponential in (2) is to numerically integrate the trajectory-marginalized equation (1). Runge-Kutta, designed for ordinary differential equations [32], balances stability and accuracy flexibly based on a chosen order of expansion and time step $\Delta t$. For example, first order Runge-Kutta reads (1)
$\bar{\rho}(t+\Delta t)\approx\bar{\rho}(t)+\Delta t\frac{\mathrm{d}\bar{\rho}}{\mathrm{d}t}=\bar{\rho}(t)+\Delta t\ \bar{\rho}(t)\mathbf{A}_{\theta}\ .$ (17)
Thus, we evolve the initial vector, $\bar{\rho}(0)$, and compute successive points along the solution curve until a final time $T$, i.e., compute $\{\bar{\rho}(\Delta t),\bar{\rho}(2\Delta t),\bar{\rho}(3\Delta t),\ldots,\bar{\rho}(T)\}$ iteratively. This is equivalent to first order Taylor on the exponential at each time step. Avoiding numerical inaccuracy and instability often warrants higher-order methods. A common choice is the fourth-order Runge-Kutta method (RK4) [32].
@@
$\begin{split}\delta_{1}&=\Delta t\ \bar{\rho}(t)\mathbf{A}_{\theta}\ ,\\ \delta_{2}&=\Delta t\ \left(\bar{\rho}(t)+\frac{\delta_{1}}{2}\right)\mathbf{A}_{\theta}\ ,\\ \delta_{3}&=\Delta t\ \left(\bar{\rho}(t)+\frac{\delta_{2}}{2}\right)\mathbf{A}_{\theta}\ ,\\ \delta_{4}&=\Delta t\ \left(\bar{\rho}(t)+\delta_{3}\right)\mathbf{A}_{\theta}\ .\end{split}$ (19)
and iteratively evolve from $t=0$ towards $t=T$. This reduces the error to $\Delta t^{4}$ [32] at each step.
@@ 2.2.5 Method #5 – Krylov subspace
The Krylov subspace technique [25, 33, 34] evolves the initial probability vector $\bar{\rho}(t)$ through time by approximating the matrix exponential over a short time step $\Delta t$, $\bar{\rho}(t+\Delta t)=\bar{\rho}(t)e^{\mathbf{A}_{\theta}\Delta t}$. In keeping with the other methods, it still avoids the full matrix exponential $e^{\mathbf{A}_{\theta}\Delta t}$, by constructing a Krylov subspace at each time step, $\mathcal{K}_{\kappa}(\mathbf{A}_{\theta},\bar{\rho}(t))$, spanned by $\{\bar{\rho}(t),\bar{\rho}(t)\mathbf{A}_{\theta},\bar{\rho}(t)\mathbf{A}_{\theta}^{2},$ $\bar{\rho}(t)\mathbf{A}_{\theta}^{3},\ldots,\bar{\rho}(t)\mathbf{A}_{\theta}^{\kappa-1}\}$, where the size of the subspace, $\kappa$, is a predefined integer which serves a role similar to the approximation order in Runge-Kutta.
Although the vectors spanning the subspace are clearly defined, we require an orthonormal basis $\{\bar{q}_{1},\bar{q}_{2},\ldots,\bar{q}_{\kappa}\}$ for $\mathcal{K}_{\kappa}(\mathbf{A}_{\theta},\bar{\rho}(t))$, in order to project $\mathbf{H}$ of $\mathbf{A}_{\theta}$ into the subspace. This projection is given by
@@
At each iteration three objects are stored: $\mathbf{A}_{\theta}$, $\mathbf{Q}$, and $\mathbf{H}$. Therefore the memory cost scales as $O(\kappa^{2}+\kappa N+[\mathbf{A}_{\theta}])$. Again assuming a) and b) mentioned in the Sec. 2.1.2 are satisfied, the total time scaling and the memory requirements scale as $O(\kappa^{2}+\kappa N^{2})$ and $O(\kappa N)$, respectively. These together determine the scaling per iteration presented in Table 1.
Note that we usually use a fixed $\kappa$ and observe the scaling for $N$, as such we are most interested in $\kappa\ll N$, but we will not take that limit in scaling explicitly as of now. This happens because, in practical terms the presence of $\kappa$ leads to overhead in the total time calculation that will be relevant in our benchmarking. However one could expect that when such limit is reached the scaling would become $O(N^{2})$ for computational time and $O(N)$ for memory, similarly to the previous methods.
@@ 3 Results
With all of the methods described, we now compare them on three systems of interest we present below (Sec. 3.1). Benchmarking of the different methods is given at Sec. 3.2.
@@ 3.1 Examples
Figure 2: Cartoon representation of dynamical systems described in Sec. 3.1. Panel a) represents a birth-death process of $R$ as described in Sec. 3.1.1, panel b) represents a system that switches between two states, only one of them, $\sigma_{\text{act}}$, produces $R$, as in Sec. 3.1.2, and panel c) represents an autoregulatory system, with one state, $\sigma_{\text{act}}$, generating a product $R$ which generates a second product $P$, which subsequently plays a role in inactivating the system by moving it to a state, $\sigma_{\text{ina}}$, where $R$ is not produced, as in Sec. 3.1.3.
@@ 3.1.1 Birth-death process
In particular in molecular biology, we can model an actively transcribing DNA sequence producing one RNA [9] with rate $\beta$ that is degraded at rate $\gamma$. Data for inference on this system would be the population of $R$ at different time points across a number of identical DNA sequences.
@@
We note that $\mathbf{\Lambda}$ is sparse by construction, and each row has, at most, two non-vanishing elements. Since $\max\limits_{n}a_{nn}$ was an important quantity when discussing the methods’ scaling, it is straightforward to see from (5) that, $\max\limits_{n}|a_{nn}|=\max\limits_{n}(\beta+\gamma n)=\beta+\gamma(N-1)$, which scales as $O(N)$.
In our synthetic experiments for this process, we fix the time units so the ground truth degradation rate is one $\gamma=1$ or, equivalently, time has units of $1/\gamma$. Crucially, this is merely a choice of parameters for the synthetic experiment and we do not assume any knowledge of $\gamma$ a priori. All simulations start with zero population of $R$, and we observe the final state at various time points $T^{i}$, evenly spaced from $0.5$ to $5$. For each of these time points, we conduct 300 individual simulations, yielding a total dataset of $I=3000$ observations, We perform synthetic experiments for each value of the production rate $\beta$ from the set $\{50,100,200,400,800\}$, selected because the population at equilibrium fluctuates around $\beta/\gamma$ allowing us to assess the scaling of each algorithm with the number of states necessary for accurate inference, spanning a range from hundreds to thousands of states.
@@ 3.1.2 Two state birth-death
Since we ought to write our state space with a single index $n$, we now index the states using $n=n_{R}+\sigma_{G}N_{R}$ with $N_{R}$ the upper cut-off for the $R$ population, and $N=2N_{R}$. Similarly, we can invert this indexing recovering the population of $R$, $n_{R}$, and the gene state $\sigma_{G}$ as $(n_{R},\sigma_{G})=\left(n\mod{N_{R}},\left\lfloor\frac{n}{N_{R}}\right\rfloor\right)$, where $x\mod q$ represents the remainder obtained when $x$ is divided by $q$ and $\lfloor x\rfloor$ represents the largest integer smaller than $x$. Here, the transition matrix elements are given by
@@
Just as we did for the previous system, we set $\gamma=1$ in our synthetic experiment thereby setting the units. In this cases, the production rate $\beta$ takes the values in $\{100,200,400,500\}$ enabling us to observe the scaling of the methods with the number of states, since the steady-state amount of product for the previous example, $\beta/\gamma$, is observed in the limiting case where the system never moves from the active state (i.e., if $\lambda_{\text{ina}}=0$). The activation and deactivation rates, $\lambda_{\text{act}}$ and $\lambda_{\text{ina}}$, are chosen to be on the same order as the measurement time, $\lambda_{\text{act}}=2$ and $\lambda_{\text{ina}}=1$. The observation points, $T^{i}$, are evenly spaced from $0.5$ to $10$, leading to a total of $I=6000$ observations. The observations (data) are the population of $R$, $n_{R}$, at the different time end points. However unlike in the previous example there is no direct observation of the $\sigma_{\text{act}}$ or $\sigma_{\text{ina}}$ state. Hence, when calculating the likelihood of observation one has to sum the probabilities of states $(n_{R},0)$ and $(n_{R},1)$.
@@ 3.2 Benchmarking
To evaluate the efficiency of the Methods (#2–#5) in avoiding the matrix exponential for a variety of physically motivated generator matrices, we embed the matrix exponential as the key computational step in a Bayesian inference scheme. To be precise, we generate synthetic data $\bar{s}$ and perform inference by sampling kinetic paramaters, $\theta$ from the posterior $p(\theta|\bar{s})$ with MCMC. The sampling scheme (see Supplemental Information A for details) is of little relevance to the matrix exponential, but we’ve made the source code publicly available on our GitHub [39].
@@
Benchmarking each method serves two primary objectives: first, it aims to evaluate time scaling with total number of states for the benchmarked methods to compare them with the conventional matrix exponential [40]. Since the conventional matrix exponential dramatically more costly than the other methods, even for this simple system, we only include it in our benchmark for the birth-death process. These results are presented and discussed in Sec. 3.2.1.
@@
Second, to ensure that each method results in an equivalent modeling scheme, we present a comparison between the probability distributions over $\theta$ obtained through various methods in the Supplemental Information B. Notably, even though the time needed to generate the samples which make up this distribution with the J-MJP (Method #2) is not considerably worse than the other methods, it generates samples that are highly correlated, a problem whose implications we expand upon in Sec. 3.2.2.
3.2.1 Time comparison

### 06 Feb 2024 v3

@@ 2.2.1 Method #1 – Trajectory-based MJP method (T-MJP)
When enough alternating samples are drawn, we obtain a model of the dynamics (that is, samples of $\theta$ informed by $\bar{s}$) avoiding entirely the matrix exponential itself.
@@ 2.2.2 Method #2 – Jump-based MJP method (J-MJP)
As with the previous method, for a sufficiently large number of alternate samples, we can avoid matrix exponentiation entirely and still obtain the appropriate probabilities for the states given the kinetic parameters, $\theta$, and the number of virtual jumps for respective observations, $\bar{k}$. In Sec. 3.2.2 we will discuss practical challenges this poses.
@@
Rather than the complete set of $\bar{\rho}(0)\mathbf{B}_{\theta}^{k}$ required by T-MJP, algorithm 1 only requires saving the vector $\bar{\rho}\mathbf{B}_{\theta}^{k}$, denoted in the pseudocode as $\bar{\rho}^{*}$. Its outer loop is run $\max\limits_{i}k^{i}$ times. Given that $k^{i}$ is Poisson distributed at rate $T^{i}\Omega_{\theta}$ (see Sec. 2.1.1) and $\Omega_{\theta}$ scales with $\max\limits_{n}{|a_{nn}|}$, the number of outer loop calls scales with $O(T\max\limits_{n}{|a_{nn}|})$, with $T$ being the largest collection time $T=\max\limits_{i}T^{i}$. Meanwhile, the most computationally intensive operation within each loop iteration is multiplying $\bar{\rho}$ by $\mathbf{B}_{\theta}$, which scales in time as $O([\mathbf{B}_{\theta}])=O([\mathbf{A}_{\theta}])$. Therefore the total time scaling of the state probability computation is $O(T\max\limits_{n}{|a_{nn}|}[\mathbf{A}_{\theta}])$. Under conditions a) and b) mentioned in the introduction, this is equivalent to $O(N^{2})$.
@@ 2.2.3 Method #3 – Remarginalized MJP method (R-MJP)
where $\varepsilon_{k^{\ast}}=1-\sum_{0}^{k^{\ast}}\ \text{Poisson}(k|\Omega_{\theta}t)$, is the upper bound in the approximation error. In practice the goal is to calculate the probability vectors at the collection times $T^{i}$. For practical reasons, we choose $k^{\ast}=\Omega_{\theta}t+6\sqrt{\Omega_{\theta}t}$, guaranteeing a modest error bound $\varepsilon_{k^{\ast}}<10^{-5}$ for $\Omega_{\theta}t>10$. While implementing the approximation in (16) is straightforward, we provide pseudocode detailing how to efficiently calculate the summation in (16) in Algorithm 2.
@@
In order to calculate the probability vector at the observation time $\bar{\rho}(t)$ from $\bar{\rho}(0)$, the outer loop in Algorithm 2 is set to run $k^{\ast}=\Omega_{\theta}t+6\sqrt{\Omega_{\theta}t}$ times. Similarly to the previous method, this scales as $O(T\max_{n}|a_{nn}|)$. Within this loop, the most computationally expensive task is multiplication of an array of $N$ elements by $\mathbf{B}_{\theta}$, with scales with $O([\mathbf{B}_{\theta}])$. Consequently, the overall time cost scaling is $O(\max_{n}|a_{nn}|[\mathbf{B}_{\theta}])$. Again, when conditions a) and b) in Sec. 2.1.2 are met, this overall cost becomes $O(N^{2})$, while the Algorithm 2 requires only saving a single array of size $N$ besides the matrix $\mathbf{B}_{\theta}$, giving memory scaling $O(N)$.
@@ 2.2.4 Method #4 – Runge-Kutta
The most conceptually straightforward technique to avoid the matrix exponential in (2) is to numerically integrate the trajectory-marginalized equation (1). Runge-Kutta, designed for ordinary differential equations [32], balances stability and accuracy flexibly based on a chosen order of expansion and time step $\Delta t$. For example, first order Runge-Kutta reads
$\bar{\rho}(t+\Delta t)\approx\bar{\rho}(t)+\Delta t\frac{\mathrm{d}\bar{\rho}}{\mathrm{d}t}=\bar{\rho}(t)+\Delta t\ \bar{\rho}(t)\mathbf{A}_{\theta}(t)\ .$ (17)
Thus, we evolve the initial vector, $\bar{\rho}(0)$, and compute successive points along the solution curve until a final time $T$, i.e., compute $\{\bar{\rho}(\Delta t),\bar{\rho}(2\Delta t),\bar{\rho}(3\Delta t),\ldots,\bar{\rho}(T)\}$ iteratively. This is equivalent to first order Taylor on the exponential at each time step.
Avoiding numerical inaccuracy and instability often warrants higher-order methods. A common choice is the fourth-order Runge-Kutta method (RK4) [32].
@@
$\begin{split}\delta_{1}&=\Delta t\ \bar{\rho}(t)\mathbf{A}_{\theta}(t)\ ,\\ \delta_{2}&=\Delta t\ \left(\bar{\rho}(t)+\frac{\delta_{1}}{2}\right)\mathbf{A}_{\theta}(t+\nicefrac{{\Delta t}}{{2}})\ ,\\ \delta_{3}&=\Delta t\ \left(\bar{\rho}(t)+\frac{\delta_{2}}{2}\right)\mathbf{A}_{\theta}(t+\nicefrac{{\Delta t}}{{2}})\ ,\\ \delta_{4}&=\Delta t\ \left(\bar{\rho}(t)+\delta_{3}\right)\mathbf{A}_{\theta}(t+{\Delta t})\ .\end{split}$ (19)
and iteratively evolve from $t=0$ towards $t=T$. This reduces the error to $\Delta t^{4}$ (see Ref.[32]) at each step. We made the time dependence in (19) explicit (writing $\mathbf{A}_{\theta}(t)$) because this method can be used for transition rates that change in time, in opposition to the previous methods relying on the expansion in (7) that assumes $\mathbf{A}_{\theta}$ is constant in time.
@@ 2.2.5 Method #5 – Krylov subspace
The Krylov subspace technique [25, 33, 34] evolves the initial probability vector $\bar{\rho}(t)$ through time by approximating the matrix exponential over a short time step $\Delta t$, $\bar{\rho}(t+\Delta t)=\bar{\rho}(t)e^{\mathbf{A}_{\theta}\Delta t}$. In keeping with the other methods, it still avoids the full matrix exponential $e^{\mathbf{A}_{\theta}\Delta t}$. It does so by constructing a Krylov subspace at each time step, $\mathcal{K}_{\kappa}(\mathbf{A}_{\theta},\bar{\rho}(t))$, spanned by $\{\bar{\rho}(t),\bar{\rho}(t)\mathbf{A}_{\theta},\bar{\rho}(t)\mathbf{A}_{\theta}^{2},$ $\bar{\rho}(t)\mathbf{A}_{\theta}^{3},\ldots,\bar{\rho}(t)\mathbf{A}_{\theta}^{\kappa-1}\}$, where the size of the subspace, $\kappa$, is a predefined integer which serves a role similar to the approximation order in Runge-Kutta. Note that, as with Methods #1 – #3, the Krylov subspace requires that $\mathbf{A}_{\theta}$ is constant in the interval $(t,t+\Delta t)$.
Although the vectors spanning the subspace are clearly defined, we require an orthonormal basis $\{\bar{q}_{1},\bar{q}_{2},\ldots,\bar{q}_{\kappa}\}$ for $\mathcal{K}_{\kappa}(\mathbf{A}_{\theta},\bar{\rho}(t))$, in order to project $\mathbf{A}_{\theta}$ into the subspace. This projection is given by
@@
At each iteration three objects are stored: $\mathbf{A}_{\theta}$, $\mathbf{Q}$, and $\mathbf{H}$. Therefore the memory cost scales as $O(\kappa^{2}+\kappa N+[\mathbf{A}_{\theta}])$. Again assuming a) and b) mentioned in the Sec. 2.1.2 are satisfied, the total time and the memory requirements scale as $O(\kappa^{2}+\kappa N^{2})$ and $O(\kappa N)$, respectively. These together determine the scaling per iteration presented in Table 1.
Note that we usually use a fixed $\kappa$ and observe the scaling for $N$, as such we are most interested in $\kappa\ll N$, but we will not take that limit in scaling explicitly as of now. This happens because, in practical terms, the presence of $\kappa$ leads to overhead in the total time calculation that will be relevant in our benchmarking. However one could expect that when such limit is reached the scaling would become $O(N^{2})$ for computational time and $O(N)$ for memory, similarly to the previous methods.
@@ 3 Results
With all of the methods described, we now compare them on three systems of interest we present below (Sec. 3.1). Benchmarking of the different methods is given in Sec. 3.2.
@@ 3.1 Examples
Figure 2: Cartoon representation of dynamical systems described in Sec. 3.1. Panel a) represents a birth-death process of $R$ as described in Sec. 3.1.1, panel b) represents a system that switches between two states, only one of them, the active one $\sigma_{\text{act}}$, produces $R$, as in Sec. 3.1.2, and panel c) represents an autoregulatory system, with one state, $\sigma_{\text{act}}$, generating a product $R$ which generates a second product $P$, which subsequently plays a role in inactivating the system by moving it to a state, $\sigma_{\text{ina}}$, where $R$ is not produced, as in Sec. 3.1.3.
@@ 3.1.1 Birth-death process
For example in molecular biology, we can model an actively transcribing DNA sequence producing one RNA [9] with rate $\beta$ degraded at rate $\gamma$. Data for inference on this system would be the population of $R$ at different time points across a number of identical DNA sequences.
@@
We note that $\mathbf{\Lambda}$ is sparse by construction, and each row has, at most, two non-vanishing elements. Since $\max\limits_{n}a_{nn}$ was an important quantity when discussing the methods’ time scaling, it is straightforward to see from (5) that, $\max\limits_{n}|a_{nn}|=\max\limits_{n}(\beta+\gamma n)=\beta+\gamma(N-1)$, which scales as $O(N)$.
In our synthetic experiments for this process, we fix the time units so the ground truth degradation rate is one $\gamma=1$ or, equivalently, time has units of $1/\gamma$. Crucially, this is merely a choice of parameters for the synthetic experiment and we do not assume any knowledge of $\gamma$ a priori. All simulations start with zero population of $R$, and we observe the final state at various time points $T^{i}$, evenly spaced from $0.5$ to $5$. For each of these time points, we conduct 300 individual simulations, yielding a total dataset of $I=3000$ observations. We perform synthetic experiments for each value of the production rate $\beta$ from the set $\{50,100,200,400,800\}$, selected because the population at equilibrium fluctuates around $\beta/\gamma$ allowing us to assess the scaling (on time and memory) of each algorithm with the number of states necessary for accurate inference, spanning a range from hundreds to thousands of states.
@@ 3.1.2 Two state birth-death
Since we must choose a linear ordering of states, in order to write their probabilities as vector, we ought to write our state space with a single index $n$. We index the states using $n=n_{R}+\sigma_{G}N_{R}$ with $N_{R}$ the upper cut-off for the $R$ population, and $N=2N_{R}$. Similarly, we can recover the population of $R$, $n_{R}$, and the gene state $\sigma_{G}$ from $n$ as $(n_{R},\sigma_{G})=\left(n\mod{N_{R}},\left\lfloor\frac{n}{N_{R}}\right\rfloor\right)$, where $x\mod q$ represents the remainder obtained when $x$ is divided by $q$ and $\lfloor x\rfloor$ represents the largest integer smaller than $x$. Here, the transition matrix elements are given by
@@
Just as we did for the previous system, we set $\gamma=1$ in our synthetic experiment thereby setting the units. In this cases, the production rate $\beta$ takes the values in $\{100,200,400,500\}$ enabling us to observe how the time and memory scales within each of the methods with the number of states, since the steady-state amount of product for the previous example, $\beta/\gamma$, is observed in the limiting case where the system never moves from the active state (i.e., if $\lambda_{\text{ina}}=0$). The activation and deactivation rates, $\lambda_{\text{act}}$ and $\lambda_{\text{ina}}$, are chosen to be on the same order as the measurement time, $\lambda_{\text{act}}=2$ and $\lambda_{\text{ina}}=1$. The observation points, $T^{i}$, are evenly spaced from $0.5$ to $10$, leading to a total of $I=6000$ observations. The observations (data) are the population of $R$, $n_{R}$, at the different time end points. However unlike in the previous example there is no direct observation of the $\sigma_{\text{act}}$ or $\sigma_{\text{ina}}$ state. Hence, when calculating the likelihood of observation one has to sum the probabilities of states $(n_{R},0)$ and $(n_{R},1)$.
@@ 3.2 Benchmarking
To evaluate the efficiency of the Methods (#2–#5) in avoiding the matrix exponential for a variety of physically motivated generator matrices, we embed the matrix exponential as the key computational step in a Bayesian inference scheme. To be precise, we generate synthetic data $\bar{s}$ and perform inference by sampling kinetic parameters, $\theta$ from the posterior $p(\theta|\bar{s})$ with MCMC. The sampling scheme (see Supplemental Information A for details) is of little relevance to the matrix exponential, but we’ve made the source code publicly available on our GitHub [36].
@@
Benchmarking each method serves two primary objectives: first, it aims to evaluate time scaling with total number of states for the benchmarked methods to compare them with the conventional matrix exponential [41]. Since the conventional matrix exponential is dramatically more costly than other methods, even for this simple system, we only include it in our benchmark for the birth-death process. These results are presented and discussed in Sec. 3.2.1.
@@
Second, to ensure that each method results in an equivalent modeling scheme, we present a comparison between the probability distributions over $\theta$ obtained through various methods in the Supplemental Information B. Notably, even though the time needed to generate the samples which make up this distribution with the J-MJP (Method #2) is not considerably worse than the other methods, it generates highly correlated samples, a problem whose implications we expand upon in Sec. 3.2.2.
3.2.1 Time scaling comparison